OpenStudy (anonymous):

Vivian is considering two different cell phone plans. If she uses an average of 500 minutes a month, which plan should she choose? True Sound is offering a cellular phone plan that costs $25 per month for 300 minutes. Any minutes after 300 will cost $0.12 a minute. Cellular Zone is offering a cellular phone plan that costs $40 per month for 400 minutes. Any minutes after 400 will cost $0.08 a minute. (1)True Sound (2)Cellular Zone (3)Either-Her cost will be the same with either plan. (4)Not enough information

OpenStudy (anonymous):

Cellular Zone total will be $40.80 where True Sound is $49

OpenStudy (anonymous):

True sound : 25 + .12(200) = 25 + 24 = 49 per month Cellular Zone : 40 + .08(100) = 40 + 8 = 48 per month If Vivian uses 500 minutes per month, it would be cheaper for her to use Cellular Zone

OpenStudy (anonymous):

is the answer 1 or 2

OpenStudy (anonymous):

Sorry, multiplied wrong haha, to many distractions Answer 2

OpenStudy (anonymous):

answer is 2. Cellular Zone

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

you get a medal

OpenStudy (anonymous):

:) happy happy happy

OpenStudy (anonymous):

Haha, no problem.